Home  Contact  Equipment  Field Notes  Photos  Data  Logs Today's Date:   
   
   
  Biomet public Home
Latest Entries
 TZ 2018-04-13 - JV
 VR 2018-04-13 - JV
 MB 2018-04-05 - DS
 WP 2018-04-05 - DS
 TA 2018-04-05 - DS
 EE 2018-04-05 - DS
 EP 2018-04-17 - DS
 SW 2018-04-05 - DS
 BA 2018-04-05 - DS
 BC 2018-04-05 - DS

Recently Added
 EP 2018-04-17 - DS
 TZ 2018-04-13 - SM
 TZ 2018-04-13 - JV
 VR 2018-04-13 - JV
 MB 2017-08-16 - JV

   
   
   
   
   
 
 
<-- | -->

East Pond: 2018-04-17 10:00:21
Author: Daphne Szutu
Others: Joe, Amy

Summary: Setting up new tower for water-level manipulation experiment in East Pilot Pond

2018-04-18
East Pond

We arrived around 10:00. It was foggy and cold, but by noon the fog had burned off and it was sunny, cool, and calm. We stopped at Home Depot on the way to the site to pick up fence posts, concrete blocks, 14/2 power cable in conduit, and pieces for another float (2” PVC 90deg street elbows). We also picked up all the scaffolding pieces and a chain at West Pond.

The tower is located on the east end of East Pond. There are tall, green tule on the edges on the pond, but the tule and cattails are much sparser and shorter (<1m tall) in the pond interior. The water is covered by the same plant as at the Open Water temporary tower last summer, the red and green plant that looks like dense moss. The ground is soft and very bubbly. As Amy walked around, she measured an average water depth of 70-80cm. We installed a water-level (staff gauge) meter stick at the base of the tower, which read 80cm.

Joe set up our solar system on one level of scaffolding about 15m north of the tower: 3 solar panels and 6 batteries (sn BB-EP) connected in a 12V system, connected to two 15A charge controllers. Some of the wires on the panels were not long enough, so the box with the charge controllers is not mounted properly. A security bolt on one plane would not tighten and should be replaced. The batteries are from the Mayberry solar system and are kind of old but tested well in the lab.

Amy set up three thermocouple profiles. The soil thermocouples are attached to a fiberglass rod and are at depths below soil surface: 2cm, 8cm, 16cm. The water themocouples are attached to a floating profile (PVC float with brass chain) at depths below water surface: 2cm, 10cm, 25cm, 40cm.
-A profile: Water was 70cm deep, about 3.5m from the tower at 275deg (west). There is a SHF plate near this profile.
-B profile: Water was 70cm deep, about 7.5-8m from the tower at 250deg (west). There is a SHF plate and a VWC sensor near this profile.
-C profile: Water was 75cm deep, about 7m from the tower at 200deg (south). There is a SHF plate and a VWC sensor near this profile.

Amy also installed a well near the southwest corner of the tower and lowered a water level sensor to the bottom of the well. Joe hung a miniDOT sensor off the west end of the tower; we need a miniUSB cable to read the data off of this sensor (no SD card).

I set up the flux, radiation, and met sensors. We dropped the HMP60 sensor in the water, and even though the sensor looked dry, the sensor was giving bad numbers, so we need to swap this next time.

I updated the 7500 configuration and pressure coefficients. I tried to change the 7550 time from 15:11 to 16:00, but I couldn’t figure out how to save the time change, so we should swap the 7550 next time. I tested the 7700 wash spin/pump; Joe adjusted the water nozzle.
The 7500 read: 411ppm CO2, 366mmol/m3 H2O, 20.4C, 102kPa, 97SS
The 7700 read: 2.0ppm CH4, 70RSSI

LI-7550: sn AIU-0666
LI-7500: sn 0418
LI-7700: sn TG1-0211
Sonic: sn 153803
HMP60: sn G4540021
Barometer (PTB110): sn H0160093
HFP (A profile): 10809/ 000137/ 002211?
HFP (B profile):
HFP (C profile):
VWC (B profile): 17131/ 17130?
VWC (C profile):
CR1000:
AM16/32:
AM25T:
Water level sensor:
MiniDOT: sn 931528
PAR incoming:
PAR outgoing:
CNR1: sn 990181

The tower still needs:
-LI-7550 (swap)
-HMP45 + shield/U-bolt + 12V fan (swap)
-eosGP
-water level tape to measure well level
-bird spikes for sonic
-deep hat for 7700
-desiccant for water level sensor
-putty/mothballs for enclosures
-conduit for thermocouples
-NDVI sensors
-small washers to lower outgoing PAR sensor
-record sensor heights
-check level of tower

2 photos found


20180417EP_towersetup.JPG ( 2018-04-17 13:14:43 ) Full size: 1080x1440
Joe installing the radiation boom on the new EP tower. The tule are tall by the edges of the levee road, but the vegetation are much sparser as you move 1-2 m into the water.


20180417EP_waterprofC.JPG ( 2018-04-17 13:18:19 ) Full size: 1920x1440
Amy installing thermocouple profile C at the East Pond site. The tule and cattail are sparse on this edge of the pond, and the open water is covered by a plant that looks like dense moss.

Graphs display: days
0 sets found

 
 

Login
Select Notes:

Search Title for:

Search Descritption for:


Equip hist
Clean html
This material is based upon work supported by the National Science Foundation and US Department of Energy. Any opinions, findings, conclusions, or recommendations expressed in the material are those of the author(s) and do not necessarily reflect the views of the supporters.